# Improving Multilingual Capabilities with Cultural and Local Knowledge in Large Language Models While Enhancing Native Performance

### v2

@@ Abstract
Large Language Models (LLMs) have shown remarkable capabilities, but their development has primarily focused on English and other high-resource languages, leaving many languages underserved. We present our latest Hindi-English bi-lingual LLM with  3% average improvement in benchmark scores over both languages, outperforming models twice its size. Using a curated dataset composed of English and Hindi instruction data of 485K samples, we instruction tuned models such as Qwen-2.5-14B-Instruct and Phi-4 to improve performance over both English and Hindi. Our experiments encompassing seven different LLMs of varying parameter sizes and over 140 training attempts with varying English-Hindi training data ratios demonstrated that it is possible to significantly improve multilingual performance without compromising native performance. Further, our approach avoids resource-intensive techniques like vocabulary expansion or architectural modifications, thus keeping the model size small. Our results indicate that modest fine-tuning with culturally and locally informed data can bridge performance gaps without incurring significant computational overhead. We release our training code, datasets, and models under mit and apache licenses to aid further research towards under-represented and low-resource languages.
@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) has led to great advances in various natural language processing tasks. However, the majority of research efforts have disproportionately focused on English and a select few high-resource languages. This disparity leaves a vast number of languages under-served, limiting the global accessibility and applicability of LLM technology. While the lack of readily available data for many languages is a contributing factor, it is not the sole reason. Economic factors and limited access to computational resources also play significant roles in accessibility to target audience. In this work, we address the gap by developing a bilingual LLM that performs well on English and Hindi tasks. We focused on maintaining relatively smaller model sizes and rather than resorting to resource-intensive methods such as vocabulary expansion, block expansion, or additional layers, we employ computationally efficient fine-tuning methods such as Supervised Fine-Tuning (SFT) (Face, 2025)(von Werra et al., 2020) with Low-Rank Adaptation (LoRA) (Hu et al., 2021) through Unsloth (Daniel Han and team, 2023). Our primary goal was to boost performance over Hindi tasks while retaining similar performance over English.
We demonstrate our method by fine-tuning Qwen-2.5-14B-Instruct (Qwen et al., 2025) and Phi-4 (Abdin et al., 2024) models on a mixed-language dataset. Moreover, our experiments extend to five other LLMs : Gemma 2 9B, Gemma 2 2B (Team, 2024a), Llama 3.1 8B, Llama 3.1 3B (Team, 2024b), Qwen 2.5 3B where over 140 fine-tuning attempts were conducted by varying the distribution ratios of Hindi and English samples of each domain in the training data. These experiments provide insights into how performance changes with varying dataset distributions over each domain. This can help in dataset curation to effectively balance bilingual performance. The promising results suggest that enhancing low-resource language capabilities doesn’t necessarily require large-scale architectural changes but can be achieved through targeted, efficient fine-tuning of models with basic capabilities over a language.
@@ 2 Related Works
Prior studies have attempted to address this disparity through various techniques, including vocabulary expansion/modification (Tejaswi et al., 2024) (Csaki et al., 2023) (Shi et al., 2024) (Balachandran, 2023), modifications in architecture like block expansion and the addition of extra layers to accommodate linguistic diversity (Llama-Nanda, 2024), or continued pre-training followed by instruction tuning again (Mahdizadeh Sani et al., 2025) (Kuulmets et al., 2024) (Cui et al., 2023) (Vo et al., 2024) (Luukkonen et al., 2023) (Kallappa et al., 2025) (Toraman, 2024). However, such methods often incur substantial computational costs and lead to increase in model sizes. Prior works also include multi-lingual LLMs optimized for several languages including Hindi : Bloom-176B (BigScienceWorkshop, 2023), Aya-23B (Aryabumi et al., 2024), Aya-101 (Üstün et al., 2024) and Aya-expanse (Dang et al., 2024). Additionally we also have several other mono-lingual and bi-lingual LLMs focused on Hindi : llama-nanda-10B (Llama-Nanda, 2024), Airavata-7B (Gala et al., 2024), (BhabhaAI, 2024), Aryabhatta-8.5B (GenVRadmin, 2024), Sarvam-2B (Sarvamai, 2024), Krutrim-2-12B (Kallappa et al., 2025) and Nemotron-mini-Hindi (Joshi et al., 2024). The key differences can be seen in Table 1.
Prior Work Model VRAM Req Approach Status
@@ 3 Datasets
Despite the existence of datasets to cover several domains for Hindi (Khan et al., 2024), (Ramesh et al., 2022), we decided to experiment primarily with translated / reformatted datasets which do not prohibit usage for research/commercial purpose. This was done so that the same work can be implemented/extended to low-resource languages. Also, fine-tuning on translated data is an efficient way to adapt mPLMs to new languages, leveraging their pre-trained multilingual knowledge. (Chen and Chen, 2024). For translation, we used GPT-4o-mini (OpenAI, 2024) through Microsoft Azure  [^1] https://azure.microsoft.com/en-us/products/ai-services/openai-service/ to translate few datasets and benchmarks from English to Hindi : Big-Bench-Hard (Suzgun et al., 2022), XNLI (Conneau et al., 2018), Xl-Sum (Hasan et al., 2021). Some of the benchmarks which already have Hindi subsets were used directly : Global MMLU (Singh et al., 2024a), IndicXNLI (Aggarwal et al., 2022). Some of the publicly available datasets containing cultural and localized general knowledge like Indian legal FAQ (Aditya2411, 2024), UPSC FAQ (prnv19, 2024), IndianTAX FAQ (msinankhan1, 2024), IndianMedicines , IndiaCuisines and IndiaTravel Guide (cyberblip, 2024) were used to generate instruction-response pairs from the tabular format data using GPT-4o-mini as a part of our dataset collection. These were first translated to the other language from the original language then manually verified by multiple annotators to ensure quality in both languages. We also used a few subsets from the Aya collection (Singh et al., 2024b) i.e the translation, simplification and summarization subsets. In total the collected dataset had 3.12M samples with nearly 50:50 ratio of English and Hindi data. Around 90K samples from these cover localized and cultural knowledge. Among the rest, some domains and tasks had higher proportion in the collection. We used randomly selected subsets from those datasets while maintaining equal language ratios. After filtering the training data, we had around 485K samples of which 20% are of localized domain and cultural knowledge, while the rest are of generic tasks like math, MCQs, reasoning, summarization, rephrasing and translation.
@@ 4 Instruction Data Formatting
During Training we have appended the inputs with different strings based on the task at hand. The details of the appended strings for each task type can be seen in Table 2. The underlined portions were replaced with the corresponding texts for each sample. This modification helped in tuning the model to obey instructions well with less additional tokens needed for formatting instructions, while not compromising the performance on both the languages. The inputs were preprocessed to replace consecutive spaces with a single space, removal of leading and trailing spaces and replacement of double quotes with single quotes. Same chat templates were used as the original models with input portions processed into our format.
Task Input Format
@@
Benchmarks Ratio of ARC-Challenge ARC-Easy MMLU BoolQ Context-MCQ Overall Average
Domain data used? Hindi En Hi En Hi En Hi En Hi En Hi En Hi Tot
@@
Benchmarks Ratio of ARC-Challenge ARC-Easy MMLU BoolQ Context-MCQ Overall Average
Domain data used? Hindi En Hi En Hi En Hi En Hi En Hi En Hi Tot
@@ 5 Initial Evaluation
Before proceeding to train over the full dataset, we have first experimented through several attempts by training on a subset of our data with/without including training data of benchmarks’ domains and by varying ratio of each language in the dataset used. The subsets contain at most 2000 samples from each dataset source for both languages combined. We used normalized next-token log probabilities for MCQs and Boolean benchmarks during the initial evaluation stage to evaluate the models. We then compared how the scores changed with these variations and compared with the original models to gather insights into optimal final dataset sampling approaches. The results over Qwen-2.5-14B and Phi-4 can be seen below in Table 3 and Table 4 respectively. The results for the rest of the models can be found in Appendix C.
@@ 6 Dataset Distribution and Ordering
The performance of models from initial tests didn’t vary significantly with/without being trained on math data. The performance on Math subsets of MMLU as well remained similar on both languages with/without being trained on math samples. Since we would be training on a large number of samples, we decided to still use a considerable amount of math samples. A significant performance gap was observed over boolean benchmarks with a nearly 3% increase in English and 5% increase in Hindi. Hence, we decided to use a slightly higher amount of boolean questions’ samples in the final dataset. The language ratios for each domain in the final dataset were determined based on the initial training data ratios that gave the best results. The samples of the final dataset were sorted over input lengths in ascending order with a certain number of longest samples placed in the beginning, this approach could improve batch processing efficiency and training stability (Wang et al., 2024a). This number was set equal to the total effective batch size (i.e the product of batch size and gradient accumulation steps). The samples related to local and cultural knowledge were then placed such that they are evenly spread out in the dataset except the initial batch. More info on the dataset can be found in Appendix B. The training methods and details can be found in Appendix A.
@@ 7 End Evaluation
Apart from the benchmarks seen in Table 3 and Table 4, we perform evaluations over additional benchmarks like : MMLU-Pro (Wang et al., 2024b), BigBench-Hard (Suzgun et al., 2022), MuSR (Sprague et al., 2024), GPQA (Rein et al., 2023), MATH-Hard (Hendrycks et al., 2021). We used open-llm-leaderboard  [^2] https://huggingface.co/spaces/open-llm-leaderboard/open_llm_leaderboard (Fourrier et al., 2024) for evaluation over some of the benchmarks through eval-harness framework(Gao et al., 2021). Table 8 demonstrates The performance of our models in comparison with the original models over several benchmarks. We did observe variations in the scores from open-llm-leaderboard and the corresponding benchmark scores which were self reported for the original models. We used the scores from the leaderboard for all models over those benchmarks for reproducibility a fair comparison. The evaluation methods used can be seen in Table 5.
Benchmark Eval Criteria Eval Framework
@@ 8 Generative tasks evaluation
Scarcity of genuine and authentic multilingual benchmarks of broad range of topics has been a concern for many languages. Prior works in comparison like (Llama-Nanda, 2024) have not included generative evaluations over either languages. while (Joshi et al., 2024) utilized limited generative benchmarks using LLM-as-a-judge to score the responses, with only the MT-Bench a translation task undergoing human evaluation. Further, training on translated data to test over benchmarks translated from English defeats the purpose of building multi-lingual and multi-cultural LLMs. (Aryabumi et al., 2024) too utilizes translated benchmarks for multilingual generative task evaluation, with additional human evaluation without topic/domain restriction. We have performed human-evaluation in the same way over both languages. These results can be seen in Figure 1. We performed human evaluations though third party annotators over both languages over few of the models that achieved comparably good performance over non-english discriminative tasks. A total of 3217 comparisons were done primarily in Hindi (2097) and the rest over English (1120). For a fair comparison we utilized the default hyper-parameters of each of the models.
@@
Model $\downarrow$ ARC-C ARC-E BoolQ CMCQ MMLU Average* MMLU-Pro GPQA MuSR BBH MATH
@@
Table 6: Metrics (.2f) of our and other LLMs over several English benchmarks
@@
The best and second best for each benchmark are highlighted as bold+underlined and underlined respectively
Model $\downarrow$ ARC-C ARC-E BoolQ CMCQ MMLU Average
@@
The best and second best for each benchmark are highlighted as bold+underlined and underlined respectively
@@ 9 Comparisons
For additional comparisons, we compare the performance of our models with other Hindi bilingual LLMs and other open-source LLMs which are optimized for Hindi. Due to the large variations in number of parameters of our models and other comparable models, we compare average benchmark performance versus the model size in terms of VRAM requirement. The comparisons over English and Hindi benchmarks along side our Qwen and phi models can be seen in Table 6 and Table 7. Over the benchmarks of higher difficulty, our models have consistently outperformed models over twice their size as seen in Table 6.
Benchmark Lang Qwen-2.5- Our Qwen Change Phi-4 Our Phi-4 Change
14B-Instruct
@@
Benchmark Lang Qwen-2.5- Our Qwen Change Phi-4 Our Phi-4 Change
14B-Instruct
@@ 9.1 Domain wise Performance change
The performance of our models compared to the original versions over MMLU-pro can be seen in Table 10. The type of questions the models faced through MMLU-Pro maybe of the same domain but were of different subdomains and task types compared to those in our datasets. For example, The CS benchmarks’ questions were MCQs about various areas of computer science while our training data over CS was solely from MBPP (Austin et al., 2021) which consists of a text input and a python code as an output. Further the only source of training data we used for economics consist of TAX filing FAQs over Indian context and primarily in Hindi. Hence such domains’ data usage was mentioned as N/A. The domains which had a performance boost in our models without being in training data had questions of the form of fill-mask or text completion which were similar to the training data from Winogrande-XL (Sakaguchi et al., 2021) and PIQA (Bisk et al., 2020) spanning several domains.
@@
Model $\rightarrow$ Qwen-2.5-14B Change Phi-4 Change Training
Domain $\downarrow$ Original Ours Original Ours Data Used

### v3

@@ Abstract
Large Language Models (LLMs) have shown remarkable capabilities, but their development has primarily focused on English and other high-resource languages, leaving many languages underserved. We present our latest Hindi-English bilingual LLM with  3% average improvement in benchmark scores over both languages, outperforming models twice its size. Using a curated dataset composed of English and Hindi instruction data of 485K samples, we instruction-tuned models such as Qwen-2.5-14B-Instruct and Phi-4 to improve performance over both English and Hindi. Our experiments, encompassing seven different LLMs of varying parameter sizes and over 140 training attempts with varying English-Hindi training data ratios, demonstrated that it is possible to significantly improve multilingual performance without compromising native performance. Further, our approach avoids resource-intensive techniques like vocabulary expansion or architectural modifications, thus keeping the model size small. Our results indicate that modest fine-tuning with culturally and locally informed data can bridge performance gaps without incurring significant computational overhead. We release our training code, datasets, and models under MIT and Apache licenses to aid further research towards under-represented and low-resource languages.
@@ 1 Introduction
The rapid advancement of Large Language Models (LLMs) has led to great advances in various natural language processing tasks. However, the majority of research efforts have disproportionately focused on English and a select few high-resource languages. This disparity leaves a vast number of languages underserved, limiting the global accessibility and applicability of LLM technology. While the lack of readily available data for many languages is a contributing factor, it is not the sole reason. Economic factors and limited access to computational resources also play significant roles in accessibility to the target audience. In this work, we address the gap by developing a bilingual LLM that performs well on English and Hindi tasks. We focused on maintaining relatively smaller model sizes, and rather than resorting to resource-intensive methods such as vocabulary expansion, block expansion, or additional layers, we employ computationally efficient fine-tuning methods such as Supervised Fine-Tuning (SFT) (Face, 2025; von Werra et al., 2020) with Low-Rank Adaptation (LoRA) (Hu et al., 2021) through Unsloth (Daniel Han and team, 2023). Our primary goal was to boost performance over Hindi tasks while retaining similar performance over English.
We demonstrate our method by fine-tuning Qwen-2.5-14B-Instruct (Qwen et al., 2025) and Phi-4 (Abdin et al., 2024) models on a mixed-language dataset. Moreover, our experiments extend to five other LLMs: Gemma 2 9B, Gemma 2 2B (Team, 2024a), Llama 3.1 8B, Llama 3.1 3B (Team, 2024b), and Qwen 2.5 3B, where over 140 fine-tuning attempts were conducted by varying the distribution ratios of Hindi and English samples of each domain in the training data. These experiments provide insights into how performance changes with varying dataset distributions over each domain. This can help in dataset curation to effectively balance bilingual performance. The promising results suggest that enhancing low-resource language capabilities doesn’t necessarily require large-scale architectural changes but can be achieved through targeted, efficient fine-tuning of models with basic capabilities over a language.
@@ 2 Related Works
Prior studies have attempted to address this disparity through various techniques, including vocabulary expansion/modification (Tejaswi et al., 2024; Csaki et al., 2023; Shi et al., 2024; Balachandran, 2023), modifications in architecture like block expansion and the addition of extra layers to accommodate linguistic diversity (Llama-Nanda, 2024), or continued pre-training followed by instruction tuning again (Mahdizadeh Sani et al., 2025; Kuulmets et al., 2024; Cui et al., 2023; Vo et al., 2024; Luukkonen et al., 2023; Kallappa et al., 2025; Toraman, 2024). However, such methods often incur substantial computational costs and lead to an increase in model sizes. Prior works also include multilingual LLMs optimized for several languages, including Hindi: Bloom-176B (BigScienceWorkshop, 2023), Aya-23B (Aryabumi et al., 2024), Aya-101 (Üstün et al., 2024), and Aya-expanse (Dang et al., 2024). Additionally, we also have several other monolingual and bilingual LLMs focused on Hindi: llama-nanda-10B (Llama-Nanda, 2024), Airavata-7B (Gala et al., 2024), (BhabhaAI, 2024), Aryabhatta-8.5B (GenVRadmin, 2024), Sarvam-2B (Sarvamai, 2024), Krutrim-2-12B (Kallappa et al., 2025), and Nemotron-mini-Hindi (Joshi et al., 2024). The key differences can be seen in Table 1.
[…]
@@ 3 Datasets
Despite the existence of datasets to cover several domains for Hindi (Khan et al., 2024), (Ramesh et al., 2022), we decided to experiment primarily with translated/reformatted datasets that do not prohibit usage for research/commercial purposes. This was done so that the same work can be implemented/extended to low-resource languages. Also, fine-tuning on translated data is an efficient way to adapt mPLMs to new languages, leveraging their pre-trained multilingual knowledge. (Chen and Chen, 2024). For translation, we used GPT-4o-mini (OpenAI, 2024) through Microsoft Azure  [^1] https://azure.microsoft.com/en-us/products/ai-services/openai-service/ to translate a few datasets and benchmarks from English to Hindi: Big-Bench-Hard (Suzgun et al., 2022), XNLI (Conneau et al., 2018), and XL-Sum (Hasan et al., 2021). Some of the benchmarks that already have Hindi subsets were used directly: Global MMLU (Singh et al., 2024a), IndicXNLI (Aggarwal et al., 2022). Some of the publicly available datasets containing cultural and localized general knowledge, like Indian legal FAQ (Aditya2411, 2024), UPSC FAQ (prnv19, 2024), IndianTAX FAQ (msinankhan1, 2024), IndianMedicines, IndiaCuisines, and IndiaTravel Guide (cyberblip, 2024), were used to generate instruction-response pairs from the tabular format data using GPT-4o-mini as a part of our dataset collection. These were first translated to the other language from the original language and then manually verified by multiple annotators to ensure quality in both languages. We also used a few subsets from the Aya collection (Singh et al., 2024b), i.e., the translation, simplification, and summarization subsets. In total the collected dataset had 3.12M samples with a nearly 50:50 ratio of English and Hindi data. Around 90K samples from these cover localized and cultural knowledge. Among the rest, some domains and tasks had a higher proportion in the collection. We used randomly selected subsets from those datasets while maintaining equal language ratios. After filtering the training data, we had around 485K samples, of which 20% are of localized domain and cultural knowledge, while the rest are of generic tasks like math, MCQs, reasoning, summarization, rephrasing, and translation.
@@ 4 Instruction Data Formatting
During training we have appended the inputs with different strings based on the task at hand. The details of the appended strings for each task type can be seen in Table 2. The underlined portions were replaced with the corresponding texts for each sample. This modification helped in tuning the model to obey instructions well with fewer additional tokens needed for formatting instructions, while not compromising the performance in both languages. The inputs were preprocessed to replace consecutive spaces with a single space, remove leading and trailing spaces, and replace double quotes with single quotes. The same chat templates were used as the original models, with input portions processed into our format.
[…]
@@
[…]
@@
[…]
@@ 5 Initial Evaluation
Before proceeding to train over the full dataset, we have first experimented through several attempts by training on a subset of our data with/without including training data of benchmarks’ domains and by varying the ratio of each language in the dataset used. The subsets contain at most 2000 samples from each dataset source for both languages combined. We used normalized next-token log probabilities for MCQs and Boolean benchmarks during the initial evaluation stage to evaluate the models. We then compared how the scores changed with these variations and compared them with the original models to gather insights into optimal final dataset sampling approaches. The results over Qwen-2.5-14B and Phi-4 can be seen below in Table 3 and Table 4, respectively. It can also be seen that, in case of Qwen, the best results were obtained when ratio of Hindi is higher than 50% but for Phi-4, the results were better with ratio of Hindi less than 50%.The results for the rest of the models can be found in Appendix D.
@@ 6 Dataset Distribution and Ordering
The performance of models from initial tests didn’t vary significantly with/without being trained on math data. The performance on math subsets of MMLU as well remained similar in both languages with/without being trained on math samples. Since we would be training on a large number of samples, we decided to still use a considerable amount of math samples. A significant performance gap was observed over Boolean benchmarks with a nearly 3% increase in English and a 5% increase in Hindi. Hence, we decided to use a slightly higher amount of Boolean questions’ samples in the final dataset. The language ratios for each domain in the final dataset were determined based on the initial training data ratios that gave the best results. The samples of the final dataset were sorted over input lengths in ascending order with a certain number of the longest samples placed in the beginning; this approach could improve batch processing efficiency and training stability (Wang et al., 2024a). This number was set equal to the total effective batch size (i.e., the product of batch size and gradient accumulation steps). The samples related to local and cultural knowledge were then placed such that they are evenly spread out in the dataset except for the initial batch. More info on the dataset can be found in Appendix C. The training methods and details can be found in Appendix A.
@@ 7 End Evaluation
Apart from the benchmarks seen in Table 3 and Table 4, we perform evaluations over additional benchmarks like like like MMLU-Pro (Wang et al., 2024b), BigBench-Hard (Suzgun et al., 2022), MuSR (Sprague et al., 2024), GPQA (Rein et al., 2023), and MATH-Hard (Hendrycks et al., 2021). We used open-llm-leaderboard  [^2] https://huggingface.co/spaces/open-llm-leaderboard/open_llm_leaderboard (Fourrier et al., 2024) for evaluation over some of the benchmarks through the eval-harness framework (Gao et al., 2021). Table 8 demonstrates the performance of our models in comparison with the original models over several benchmarks. We did observe variations in the scores from the open-llm-leaderboard and the corresponding benchmark scores, which were self-reported for the original models. We used the scores from the leaderboard for all models over those benchmarks for reproducibility and a fair comparison. The evaluation methods used can be seen in Table 5.
[…]
@@ 8 Generative tasks evaluation
Scarcity of genuine and authentic multilingual benchmarks of a broad range of topics has been a concern for many languages. Prior works in comparison, like (Llama-Nanda, 2024), have not included generative evaluations over either language. while (Joshi et al., 2024) utilized limited generative benchmarks using LLM-as-a-judge to score the responses, with only the MT-Bench, a translation task, undergoing human evaluation. Further, training on translated data to test over benchmarks translated from English defeats the purpose of building multilingual and multicultural LLMs. (Aryabumi et al., 2024) also utilizes translated benchmarks for multilingual generative task evaluation, with additional human evaluation without topic/domain restriction. We have performed human evaluation in the same way over both languages. These results can be seen in Figure 1. We performed human evaluations through third-party annotators over both languages over a few of the models that achieved comparably good performance over non-English discriminative tasks. A total of 3217 comparisons were done primarily in Hindi (2097) and the rest in English (1120). For a fair comparison, we utilized the default hyperparameters of each of the models.
@@
[…]
@@
Table 6: Metrics (.2f) of our and other LLMs over several English benchmarks.
@@
[…]
@@ 9 Comparisons
For additional comparisons, we compare the performance of our models with other Hindi bilingual LLMs and other open-source LLMs that are optimized for Hindi. Due to the large variations in the number of parameters of our models and other comparable models, we compare average benchmark performance versus the model size in terms of VRAM requirement. The comparisons over English and Hindi benchmarks alongside our Qwen and Phi models can be seen in Table 6 and Table 7. Over the benchmarks of higher difficulty, our models have consistently outperformed models over twice their size, as seen in Table 6. The comparison also includes a version of Qwen that was trained on purely translated data, unlike the other two, where a translated dataset is used in cases of missing data. This was done by translating the original dataset to English and back-translating back to Hindi.  [^3] Qwen trained purely on translated data produced nearly identical results on both languages’ benchmarks compared to the main model trained on a mix of real and translated data.
[…]
@@
[…]
@@ 9.1 Domain wise Performance change
The performance of our models compared to the original versions over MMLU-pro can be seen in Table 10. The type of questions the models faced through MMLU-Pro may be of the same domain but were of different subdomains and task types compared to those in our datasets. For example, the CS benchmarks’ questions were MCQs about various topics in computer science, while our training data over CS was from MBPP (Austin et al., 2021) alone, which consists of a text input and a Python code as an output. Further, the only source of training data we used for economics consists of tax filing FAQs in the Indian context and primarily in Hindi. Hence, such domains’ data usage was mentioned as N/A. The domains that had a performance boost in our models without being in training data had questions of the form of fill-mask or text completion, which were similar to the training data from Winogrande-XL (Sakaguchi et al., 2021) and PIQA (Bisk et al., 2020) spanning several domains.
@@
[…]